niahmk:

A square pyramid is shown below: A square pyramid is shown. The sides of the square base are labeled 0.9 foot. The height of one of the triangular sides is labeled 4 feet What is the surface area of the pyramid? 8.01 square feet 2.61 square feet 4.41 square feet 1.71 square feet

niahmk:

1 attachment
AZ:

Surface area of a square pyramid You have to find the area of each side of the pyramid and add it all up We have 4 triangles and 1 square area of a triangle = 1/2 * base * height the base is 0.9 the height is 4 what is the area of one triangle?

niahmk:

3.6?

@az wrote:
Surface area of a square pyramid You have to find the area of each side of the pyramid and add it all up We have 4 triangles and 1 square area of a triangle = 1/2 * base * height the base is 0.9 the height is 4 what is the area of one triangle?

AZ:

No 1/2 * 4 * 0.9 = ?? You forgot to multiply by 1/2 or 0.5

AZ:

and remember, that's the area of just one triangle so then we have to multiply that by 4 to get the area of 4 triangles and we also still need to find the area of the square area of square = side * side and the side length is 0.9

niahmk:

@az wrote:
No 1/2 * 4 * 0.9 = ?? You forgot to multiply by 1/2 or 0.5
thats 1.8 right?

AZ:

yes and now multiply by 4 that tells you the area of the 4 triangles of the pyramid

AZ:

so what is 1.8 * 4 =